A man pulls a box with a force of 70N along the positive y axis, while a boy pulls it with a force of 50N, making an angle of 80° counter clockwise with the positive y axis. In what direction will the second boy apply a force of 60N so that the box will move in the positive y axis

1
Expert's answer
2022-04-22T13:53:31-0400

"\beta=\arctan\frac{70-50\sin 80°}{60}=19.3°."